OpenStudy (anonymous):

consider the curve x+xy+2y^2=6. the slope of the line tangent to the curve at the point (2,1) is?

OpenStudy (amistre64):

1 + xy'+y+4y y' = 0 y' = -1-y/(x+4y)

OpenStudy (amistre64):

0

OpenStudy (anonymous):

The derivative of your function is 1 + (1*dy/dx + y) - 4y(dy/dx) = 0 2 + dy/dx + y - 4y(dy/dx) = 0 (2 + y) + dy/dx (1 - 4y) = 0 (dy/dx) = -(2 + y)/(1 - 4y) and at point (2,1), dy/dx = -(2 + 1)/(1 - 4*1) = -3 / -3 = 1 Slope of tangent line = 1

OpenStudy (amistre64):

i think your derivative is off a bit... but it could be me

OpenStudy (amistre64):

x; 1 xy; xy' + y 2y^2; 4y y' = 6; 0 1 +xy' +y +4y y' = 0 xy' +4y y' = -1-y y'(x+4y) = (-1-y) y' = (-1-y)/(x+4y) ; at (x=2,y=1) y'= -1-(1)/..... = 0

OpenStudy (amistre64):

your product rule on (xy) is a little off

OpenStudy (amistre64):

and i cant seem to add lol -1-1 = -2 =-1/3 2+4 = 6